# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 5.3 Certification and Reflection
    Mechanized. Cubic compilation, degree bounds, and checker–constraint equivalence are verified in Rocq ($\mathtt{R00}$$\mathtt{R19}$). This tier also contains a code-level diagonal: the HBL conditions ($\mathtt{P01}$), fixed-point construction ($\mathtt{P02}$), and the coded undecidability theorem $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) are verified without classical logic ($\mathtt{P00}$$\mathtt{P06}$).
  2. Tier 2.
    Classical. Section 5.2 extends the result to arbitrary Turing Machines via metatheoretic LEM. This is the classical counterpart of $\mathtt{no\_total\_correct\_code\_CubicSat}$.
  3. Tier 3.
    Internal. $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ implies $\Sigma_{1}$-reflection (Theorem 5.6), contradicting Gödel’s 2nd (Corollary 5.7). This corresponds to $\mathtt{RA\_Certification}$ ($\mathtt{P04}$).
No tier conflates constructive compiler correctness with classical self-reference. ∎
@@ 6 Cubic Thresholds
6.1 On the Universal Cubic
Intuition.
The preceding construction already furnishes a uniform primitive recursive compiler from arithmetic sentences to single cubic equations, suggesting that universality at degree $3$ is not obstructed. Achieving a fixed universal cubic would require internalizing this compiler into one static cubic template while preserving strict degree control via band isolation and guard activation.
Conjecture 6.1 (Universal Cubic Conjecture).
There must be a finite equation,
$U(\vec{x},n)\in\mathbb{Z}[\vec{x},n]$ (6.125)
of total degree $\leq 3$ such that for every recursively enumerable set $W\subseteq\mathbb{N}$ there exists a parameter $n_{W}\in\mathbb{N}$ with
$k\in W\;\Longleftrightarrow\;\exists\vec{y}\in\mathbb{N}^{m}\;U(k,\vec{y},n_{W})=0.$ (6.126)
@@ 6.2 Cubic Upper Bound
Intuition.
The result obtained below is a quick and dense worst-case closure bound. It is not intrinsic to degree $3$, and one could certainly lower it by introducing only the multiplicative wiring actually used by the chosen quartic universal polynomial, as shown in Corollary 6.4.
Lemma 6.2 (Quartic-to-cubic closure).
@@
$\mathrm{H}{\mathbb{N}}(4,v)\;\leq_{m}\;\mathrm{H}{\mathbb{N}}(3,\,v+\binom{v+1}{2}\,)$. (6.127)
Introduce auxiliary variables $p_{ij}(1\leq i\leq j\leq v)$ subject to $p_{ij}=x_{i}x_{j}.$ Each such constraint has degree $2$ and every quartic monomial $x_{a}x_{b}x_{c}x_{d}$ occurring in $P$ is rewritten as $x_{a}x_{b}x_{c}x_{d}\;\longmapsto\;p_{ab}\,p_{cd},$ applying a fixed pairing rule (e.g., $a\leq b$, $c\leq d$, $ab\leq cd$ lexicographically), let $\widetilde{P}(x,p)$ denote the resulting expression. (i) Each constraint $p_{ij}=x_{i}x_{j}$ has degree $2$ in $(x,p)$. (ii) Each rewritten quartic monomial $x_{a}x_{b}x_{c}x_{d}$ is replaced by the new monomial $p_{ab}\,p_{cd}$, which is quadratic in the auxiliary variables and therefore has total syntactic degree $2$ in the enlarged variable set $(x,p)$. (iii) All cubic and lower-degree terms of $P$ are unchanged, hence have degree $\leq 3$ in $(x,p)$. Therefore, every equation in the resulting system has total degree $\leq 3$. Solvability is preserved for $\exists x\in\mathbb{N}^{v}\;P(x)=0$ if and only if,
$\exists(x,p)\in\mathbb{N}^{v+\binom{v+1}{2}}\;\Bigl(\widetilde{P}(x,p)=0\;\wedge\;\forall\,i\leq j:\;p_{ij}=x_{i}x_{j}\Bigr).$ (6.128)
Corollary 6.3 (Trivial cubic threshold).
@@
$\mathrm{H}{\mathbb{N}}(4,58)={\rm U}.$ (6.129)
Hence, by Lemma 1.5 and Corollary 5.3 it can be said,
$\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}.$ (6.130)
Follows from Lemma 6.2. The number of auxiliary variables is
$\binom{59}{2}=\frac{59\cdot 58}{2}=1711.$ (6.131)
The total variable count is therefore $58+1711=1769$. Since $\mathrm{H}{\mathbb{N}}(4,58)={\rm U}$, Lemma 6.2 yields $\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}$. ∎
Corollary 6.4 (Sparsity).
@@
Repeat the construction of Lemma 6.2, but introduce an auxiliary variable $p_{ij}$ only for pairs $(i,j)$ that actually appear in the chosen quartic-monomial decomposition of $P$. No other product variables are used in the rewritten system, so the number of added variables is exactly $m_{2}(P)$. The same back-substitution argument gives equivalence of solvability. ∎

### v7

@@ 5.3 Certification and Reflection
By mechanization. The argument has three strictly separated tiers: cubic compilation, degree bounds, and checker–constraint equivalence are verified in Rocq (core chain $\mathtt{R00}$$\mathtt{R19}$, with auxiliary support in $\mathtt{R20}$$\mathtt{R22}$). This tier also contains a code-level diagonal: the HBL conditions ($\mathtt{P01}$), fixed-point construction ($\mathtt{P02}$), and the coded undecidability theorem $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) are verified without classical logic ($\mathtt{P00}$$\mathtt{P06}$). Section 5.2 extends the result to arbitrary Turing Machines via metatheoretic LEM. This is the classical counterpart of $\mathtt{no\_total\_correct\_code\_CubicSat}$. $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ implies $\Sigma_{1}$-reflection (Theorem 5.3.4), contradicting Gödel’s 2nd (Corollary 5.3.5). This corresponds to $\mathtt{RA\_Certification}$ ($\mathtt{P04}$). No tier conflates constructive compiler correctness with classical self-reference. ∎
6 Universal Cubic
Remark 6.0.1.
This section constructs concrete and effective cubic polynomials that recover bounded proof checking for $\mathsf{RA}$ once combined with the semantic bridge of Section 5. Each artifact is a computable witness: a pinned coefficient table one can evaluate, a degree certificate checked by the Rocq kernel, and a bounded satisfiability predicate that is decidable in principle by finite search. The bounded construction is more informative than a classical existence proof: where classical universality asserts that some polynomial encodes all $\mathsf{RE}$ sets, the effective construction gives explicit coefficients, a kernel-checked degree bound, and a mechanized integrity gate.
Correctness rests on two independent layers that are jointly stronger than either alone. Rocq proves the universal statement: compilation preserves satisfiability and the aggregated polynomial has degree $\leq 3$, for every proof and every formula. The Python generator demonstrates that the claim is not vacuous: it produces live artifacts with verifiable witnesses—concrete satisfying assignments one can evaluate—whose coefficients are published, pinned by a modular digest, and kernel-checked at import time. If the mathematical reduction were unsound, the generator would fail to produce a consistent witness; it does not. Classical universality follows by taking the $\mathsf{RE}$ union over resource bounds; incompleteness manifests at the quantifier boundary where that union admits no computable collapse.
Mechanization 6.0.2 (Proof environment T003).
The concrete universal cubic is mechanized in
Repository: github.com/…/Proofcase/…/T003/
The artifact is a pinned coefficient table generated as Rocq data: monomials are stored in $20$ chunk files ($\mathtt{Coeff\_chunk\_0}$$\mathtt{Coeff\_chunk\_19}$) and assembled as $\mathtt{U\_table}$ in $\mathtt{R04\_\_Universal.v}$. The record $\mathtt{UC}$ bundles the monomial list, evaluation, the public-input index $\mathtt{ix\_u}$, and an integrity gate based on a digest of the table.
The mechanization validates syntactic and integrity properties of this pinned object (degree bound, digest agreement, basic table well-formedness) and provides the bounded satisfiability predicate used throughout this section. The semantic link to proof checking (i.e. the equivalence between bounded provability and the existence of a bounded satisfying environment) is established in a separate semantic bridge layer (Section 5).
  1. (i)
    $\mathtt{R03}$. $\mathtt{poly\_deg\_ok}$ checks that every monomial has total degree $\leq 3$. The theorem $\mathtt{degree\_le\_3}$ discharges this by $\mathtt{vm\_compute}$; $\mathtt{degree\_at\_most\_3}$ lifts the result to $\mathtt{max\_degree}\leq 3$.
  2. (ii)
    $\mathtt{R07}$. A modular digest $\mathtt{table\_digest\_of}$ is computed over the monomial list. The gate $\mathtt{table\_digest\_ok}$ certifies equality with the expected digest (generated by the Python constructor) via $\mathtt{vm\_compute}$.
  3. (iii)
    $\mathtt{R05}$. Basic well-formedness checks are provided for the pinned table: indices are within the declared variable range, no monomial has zero coefficient, the table is non-empty, and evaluation at concrete environments is computable.
  4. (iv)
    $\mathtt{R08}$. The bounded satisfiability predicate is defined as
    $\mathtt{Bounded\_Thm\_k}\;t\;:=\;\exists\,\rho\;\bigl(\mathtt{InputEncoding\_table}\;t\;\rho\\ \;\land\;\mathtt{RepresentationBound}\;\rho\;\land\;\mathtt{U\_table}\;\rho=0\bigr).$ (148)
    The lemma $\mathtt{bounded\_universal\_cubic}$ is an interface lemma identifying $\mathtt{U\_table}$ with $\mathtt{uc\_U\ UC}$ (the evaluator packaged in the record), allowing later results to refer uniformly to the record-level evaluator.
The Python generator (gen.py) supports multiple parameter profiles and emits coefficient chunks, a variable map, and metadata modules. The Rocq development in $\mathtt{T003}$ pins a specific generated table and enforces its integrity via the digest gate (table_digest_ok) together with a kernel-checked degree certificate. Table 2 reports two representative generator profiles (a minimal and an effective configuration).
Parameter Minimal ($\mathsf{RE}$) Effective ($\mathrm{imp},\bot,\mathsf{EFQ}$)
Focus $\mathsf{Rocq}$ Proof of concept
Radix base 55 6,000,0016{,}000{,}001
Proof lines ($k$) 33 11
Digit width $16$ $36$
Variables 18,71218{,}712 13,54513{,}545
Aggregation channels 6,6756{,}675 4,8424{,}842
Monomials 21,47121{,}471 15,41715{,}417
Max coefficient digits $198$ 1,2341{,}234
Table 2: Concrete bounded cubic artifacts generated by the same compiler with different parameter choices. Both are degree-33 bounded proof-check polynomials; they differ by engineering bounds and aggregation scale.
The minimal artifact supports proofs of at most $3$ lines with representation bound $\mathtt{max\_repr\_full}=2{,}583$ and lane bound $\mathtt{max\_repr\_lane}=7$; each line formula code must fit within the lane bound. The effective artifact uses a large base (6,000,0016{,}000{,}001) with $1$ proof line and wider digit width, enabling richer formula encodings within a single line. Each artifact is decidable (finite search space $(B+1)^{N}$), recursively enumerable, and able to validate bounded proofs according to its pinned parameters. Both are generated by the same Python compiler with different parameter choices; the Rocq modules are generated against the minimal profile.
6.1 Bounded Universality
Definition 6.1.1 (Universal cubic family).
For each resource parameter $k\in\mathbb{N}$, the compiler of Sections 44.9 together with mixed-radix aggregation (Lemma 4.8.5) produces a polynomial
$U_{k}(u,x_{1},\dots,x_{N_{k}})\in\mathbb{Z}[u,x_{1},\dots,x_{N_{k}}]$ (149)
of total degree $\leq 3$, given effectively by a finite coefficient table. The variable count $N_{k}$ and the aggregation base $B_{k}^{\mathrm{agg}}$ are computable functions of $k$; the constraint system $\mathcal{C}_{n,t}$ is uniform across $k$, and the only parameter that changes is the base and the resulting coefficient table. A concrete artifact (such as $\mathtt{T003}$) is a single pinned member $U_{k_{0}}$ of this family. We observe $U_{k}$ over $\mathbb{Z}$-valued environments (Definition 6.1.3) but restrict attention to nonnegative bounded assignments when defining satisfiability (Definition 6.1.6), which corresponds canonically to $\mathbb{N}$-solutions.
Remark 6.1.2 (Effectivity versus existence).
The effective construction is orthogonal to classical existence results. A classical universal Diophantine equation (as in MRDP) asserts that some polynomial exists whose solution set encodes all $\mathsf{RE}$ sets; the proof is ontological. Here, the polynomial is given: its coefficients are published (Table 2), its degree is kernel-checked ($\mathtt{degree\_le\_3}$), and its bounded universality is mechanized ($\mathtt{bounded\_universal\_cubic}$). Each bounded artifact is a constructive witness of $\mathsf{RE}$-ness for its slice. The classical result—that no uniform decider exists for the compiled family—follows indirectly: each fixed slice is decidable, yet the $\mathsf{RE}$ set $\mathrm{Sat}(t)=\exists k\,\mathrm{Sat}_{B_{k}}(t)$ resists computable collapse, since any such collapse would decide $\mathsf{RA}$-theoremhood.
Definition 6.1.3 (Environment and evaluation domain).
An environment is a function
$\rho\colon\{0,\dots,N-1\}\to\mathbb{Z}.$ (150)
We fix a distinguished index
$i_{u}\in\{0,\dots,N-1\}$ (151)
and inject the public input $t\in\mathbb{N}$ by requiring $t$ to be an integer: $\rho(i_{u})=t$.
The universal cubic $U_{k}(\rho)\in\mathbb{Z}$ is evaluated by the standard substitution semantics for integer-coefficient polynomials. Although $U_{k}$ has coefficients in $\mathbb{Z}$, all witnesses of interest are nonnegative environments: bounded satisfaction (Definition 6.1.6) explicitly enforces $\rho(i)\in[0,B]$ for all relevant indices.
Remark 6.1.4 (Why $\mathbb{Z}$-valued environments appear).
The Diophantine normal forms used in the mechanization are most convenient over $\mathbb{Z}$ because intermediate constraint languages employ subtraction-like rearrangements when compiling and simplifying equations. This does not change the Diophantine content: the satisfiability predicates quantify over environments $\rho$ that are explicitly bounded in the nonnegative range $[0,B]$, so they correspond canonically to $\mathbb{N}$-assignments by restriction. In particular, all semantic statements about bounded satisfaction are invariant under the $\mathbb{N}\hookrightarrow\mathbb{Z}$ embedding.
Definition 6.1.5 (Representation bound).
For $B\in\mathbb{N}$, define the boundedness predicate
$\mathrm{Bound}_{B}(\rho)\;:\iff\;\forall\,i<N,\;0\leq\rho(i)\leq B,$ (152)
where inequalities are taken in $\mathbb{Z}$ (and $B$ is coerced to $\mathbb{Z}$).
Definition 6.1.6 (Bounded satisfaction).
For a fixed resource parameter $k\in\mathbb{N}$ with associated bound $B_{k}$ and polynomial $U_{k}$,
$\mathrm{Sat}_{B_{k}}(t)\;:\iff\;\exists\,\rho\;\big(\rho(i_{u})=t\;\wedge\;\mathrm{Bound}_{B_{k}}(\rho)\;\wedge\;U_{k}(\rho)=0\big).$ (153)
Proposition 6.1.7 (Decidability of bounded slices).
For each fixed $k\in\mathbb{N}$, the predicate $\mathrm{Sat}_{B_{k}}(t)$ is decidable.
By exhaustion. The search space has cardinality $(B_{k}+1)^{N_{k}+1}$, which is finite for fixed $k$. Evaluation of $U_{k}$ at any integer point is computable. Hence $\mathrm{Sat}_{B_{k}}(t)$ is decided by exhaustive enumeration. ∎
The semantic claim is parameterized by a resource bound. Let $\mathsf{RA}$ be the formal system of Sections 1.13. Let $\operatorname{Prov}_{k}(t)$ denote: “there exists a proof of the formula coded by $t$ with resource bound kk” (e.g. number of lines, digit width, or trace length).
Theorem 6.1.8 (Bounded universality).
For each fixed resource parameter $k\in\mathbb{N}$, there exists a computable bound $B_{k}\in\mathbb{N}$ such that for all $t\in\mathbb{N}$,
$\operatorname{Prov}_{k}(t)\;\Longleftrightarrow\;\mathrm{Sat}_{B_{k}}(t).$ (154)
By correctness. Fix $k$. A proof with resource bound $k$ is a finite object whose encoding has size bounded by a computable function of $k$. The compiler of Sections 44.9 maps each such proof to an environment $\rho$ satisfying $U_{k}(\rho)=0$, with $\|\rho\|_{\infty}$ bounded computably in the encoding size. Set $B_{k}$ to this computable bound; the aggregation base of $U_{k}$ is chosen large enough that all channel values at any assignment in $0,B_{k}]^{N}$ lie below the base, so the capacity constraints (Definition 4.8.3) built into $U_{k}$ are satisfiable for every valid proof witness in the slice. Self-bounding aggregation (Lemma 4.8.5) then yields the equivalence. Conversely, any environment $\rho$ with $\mathrm{Bound}_{B_{k}}(\rho)$ and $U_{k}(\rho)=0$ decodes to a valid proof of resource bound $\leq k$, by the reverse direction of Lemma 4.8.5 (the capacity constraints force digit bounds, enabling extraction) and the correctness of the encoding (Theorem 1.3.1). ∎
Remark 6.1.9.
The constraint system $\mathcal{C}_{n,t}$ is independent of $k$; only the representation bound $B_{k}$ and the aggregation base change. The aggregated polynomial $U_{k}$ has the base baked into its coefficients (as powers $B^{i}$ weighting each channel). Different resource parameters $k$ may therefore require different aggregated polynomials. For a fixed artifact (such as $\mathtt{T003}$), $k$ and the base are pinned at generation time; the bounded universality theorem applies to that specific slice.
6.2 From Bounded to Classical Universality
Remark 6.2.1.
From an effective realization, classical results follow.
Lemma 6.2.2 (Existential bound extraction).
Define the unbounded satisfiability predicate:
$\mathrm{Sat}(t)\;:\iff\;\exists\,k\in\mathbb{N}\;\;\mathrm{Sat}_{B_{k}}(t),$ (155)
where $B_{k}$ and the corresponding aggregated polynomial $U_{k}$ are those of Theorem 6.1.8 and Remark 6.1.9. Equivalently,
$\mathrm{Sat}(t)\;\iff\;\exists\,k\in\mathbb{N}\;\exists\,\rho\;\big(\rho(i_{u})=t\;\wedge\;\mathrm{Bound}_{B_{k}}(\rho)\;\wedge\;U_{k}(\rho)=0\big).$ (156)
Immediate. The two formulations are equivalent by unfolding $\mathrm{Sat}_{B_{k}}$. ∎
Remark 6.2.3.
The above is elementary but decisive. It isolates the quantifier shift: unbounded provability is existential over resource bounds plus bounded search. Each slice $\mathrm{Sat}_{B_{k}}$ involves a single cubic equation $U_{k}$ with fixed base; the existential over $k$ absorbs the dependence of the base on the proof size.
Proposition 6.2.4 (Classical universality).
Let $\operatorname{Prov}(t)$ denote ordinary (unbounded) provability in $\mathsf{RA}$. Under the correctness of the encoding (Theorem 1.3.1),
$\operatorname{Prov}(t)\;\iff\;\mathrm{Sat}(t).$ (157)
By correctness.
Forward. Any proof has finite size, hence fits within some resource bound $k$. By Theorem 6.1.8, $\operatorname{Prov}_{k}(t)\implies\mathrm{Sat}_{B_{k}}(t)\implies\mathrm{Sat}(t)$.
Reverse. $\mathrm{Sat}(t)$ means $\exists k\;\mathrm{Sat}_{B_{k}}(t)$. The reverse direction of bounded universality (Theorem 6.1.8) gives $\operatorname{Prov}_{k}(t)$, hence $\operatorname{Prov}(t)$. ∎
6.3 Decidability versus Semi-Decidability
Proposition 6.3.1.
Each slice $\mathrm{Sat}_{B_{k}}(t)$ is decidable (Proposition 6.1.7).
Corollary 6.3.2.
$\mathrm{Sat}(t)$ is recursively enumerable.
By construction. Enumerate pairs $(k,\rho)$ with $\mathrm{Bound}_{B_{k}}(\rho)$ and test $U_{k}(\rho)=0$. ∎
Proposition 6.3.3 (Computable bounding implies decidability).
The critical structural observation is that the RE predicate $\mathrm{Sat}(t)$ cannot collapse to any single decidable slice. If there existed a computable function $f\colon\mathbb{N}\to\mathbb{N}$ such that
$\mathrm{Sat}(t)\;\iff\;\mathrm{Sat}_{B_{f(t)}}(t),$ (158)
then $\mathrm{Sat}(t)$ would be decidable.
Immediate. Given $t$, compute $f(t)$, then decide $\mathrm{Sat}_{B_{f(t)}}(t)$ by Proposition 6.1.7. ∎
Theorem 6.3.4 (Undecidability consequence).
If $\mathsf{RA}$ is sufficiently strong and consistent, then no computable bounding function $f$ as in Proposition 6.3.3 exists.
By reductio.
Proposition 6.2.4, $\mathrm{Sat}(t)\iff\operatorname{Prov}(t)$. The set of theorems of a consistent, sufficiently strong theory is $\mathsf{RE}$ but not decidable (Church’s theorem Church [1936]; Gödel’s first incompleteness theorem Gödel [1931]). If $f$ existed, $\mathrm{Sat}(t)$ would be decidable by Proposition 6.3.3, contradicting undecidability of $\operatorname{Prov}(t)$. ∎
6.4 Conceptual Synthesis
Remark 6.4.1.
The results of this section are summarized as follows.
  1. (i)
    Effective construction produces concrete cubic artifacts: finite coefficient tables (Table 2) with kernel-checked degree certificates and digest gates. Each artifact is a computable object, not an existence assertion.
  2. (ii)
    Bounded universality yields decidable Diophantine characterizations: for each $k\in\mathbb{N}$, the predicate $\mathrm{Sat}_{B_{k}}$ is decidable by finite search (Proposition 6.1.7). Each bounded artifact is a constructive witness of $\mathsf{RE}$-ness for its slice.
  3. (iii)
    Classical universality is their $\mathsf{RE}$ union:
    $\mathrm{Sat}\;=\;\bigcup_{k\in\mathbb{N}}\mathrm{Sat}_{B_{k}}.$ (159)
    The bounded family indirectly recovers the classical result: any $\mathsf{RE}$ set that arises as such a union is classically recognizable, and the compiler witnesses that $\mathsf{RA}$-theoremhood is among them.
  4. (iv)
    Incompleteness is the non-existence of a computable collapse from the union to a single decidable slice. No computable selector $k=f(t)$ exists (Theorem 6.3.4). This is the incompleteness phenomenon made computationally visible through the Diophantine lens.
Mechanization 6.4.2.
The contribution is an explicit, effectively generated cubic compiler tied to $\mathsf{RA}$ with mechanized bounded encoding. The $\mathtt{T003}$ artifact publishes both parameterizations (minimal and effective) as machine-readable JSON coefficient tables alongside rendered HTML equation views. The Python generator $\mathtt{gen.py}$ is the source of truth for encoding semantics; Rocq consumes the generated table and $\mathtt{VarMap}$. The universality equivalence is proved in a separate semantic bridge layer (Section 5) without normalizing the coefficient table.
[…]
Figure 3: External generation and kernel-level validation of the cubic Diophantine artifact. The Python generator produces a coefficient table (chunked) together with an expected digest; Rocq checks digest agreement and the degree-$3$ certificate, then provides evaluation and the bounded satisfiability predicate. The semantic equivalence to bounded proof checking is established in a separate bridge layer (Section 5).
@@ 7.1 Cubic Upper Bound
Remark 7.1.1.
The result obtained below is a quick and dense worst-case closure bound. It is not intrinsic to degree $3$, and one could certainly lower it by introducing only the multiplicative wiring actually used by the chosen quartic universal polynomial, as shown in Corollary 7.1.4.
Lemma 7.1.2 (Quartic-to-cubic).
@@
$\mathrm{H}{\mathbb{N}}(4,v)\;\leq_{m}\;\mathrm{H}{\mathbb{N}}(3,\,v+\binom{v+1}{2}\,)$. (160)
By construction. Introduce auxiliary variables $p_{ij}(1\leq i\leq j\leq v)$ subject to $p_{ij}=x_{i}x_{j}.$ Each such constraint has degree $2$ and every quartic monomial $x_{a}x_{b}x_{c}x_{d}$ occurring in $P$ is rewritten as $x_{a}x_{b}x_{c}x_{d}\;\longmapsto\;p_{ab}\,p_{cd},$ applying a fixed pairing rule (e.g., $a\leq b$, $c\leq d$, $ab\leq cd$ lexicographically), let $\widetilde{P}(x,p)$ denote the resulting expression. (i) Each constraint $p_{ij}=x_{i}x_{j}$ has degree $2$ in $(x,p)$. (ii) Each rewritten quartic monomial $x_{a}x_{b}x_{c}x_{d}$ is replaced by the new monomial $p_{ab}\,p_{cd}$, which is quadratic in the auxiliary variables and therefore has total syntactic degree $2$ in the enlarged variable set $(x,p)$. (iii) All cubic and lower-degree terms of $P$ are unchanged, hence have degree $\leq 3$ in $(x,p)$. Therefore, every equation in the resulting system has total degree $\leq 3$. Solvability is preserved for $\exists x\in\mathbb{N}^{v}\;P(x)=0$ if and only if,
$\exists(x,p)\in\mathbb{N}^{v+\binom{v+1}{2}}\;\Bigl(\widetilde{P}(x,p)=0\;\wedge\;\forall\,i\leq j:\;p_{ij}=x_{i}x_{j}\Bigr).$ (161)
Corollary 7.1.3 (Trivial cubic threshold).
@@
$\mathrm{H}{\mathbb{N}}(4,58)={\rm U}.$ (162)
Hence, by Lemma 1.2.2 and Corollary 5.2.3,
$\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}.$ (163)
Immediate from Lemma 7.1.2. The number of auxiliary variables is
$\binom{59}{2}=\frac{59\cdot 58}{2}=1711.$ (164)
The total variable count is therefore $58+1711=1769$. Since $\mathrm{H}{\mathbb{N}}(4,58)={\rm U}$, Lemma 7.1.2 yields $\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}$. ∎
Corollary 7.1.4 (Sparsity).
@@
By construction. Repeat the construction of Lemma 7.1.2, but introduce an auxiliary variable $p_{ij}$ only for pairs $(i,j)$ that actually appear in the chosen quartic-monomial decomposition of $P$. No other product variables are used in the rewritten system, so the number of added variables is exactly $m_{2}(P)$. The same back-substitution argument gives equivalence of solvability. ∎